# Hyperbolic recurrent neural network as the first type of non-Euclidean neural quantum state ansatz

### v2

@@ 5.3.1 Experiment details
For the 1D $J_{1}J_{2}$ system with $J_{1}=1.0$ and $J_{2}=0.0,0.2,0.5,0.8$, we carried out several VMC experiments using Euclidean and hyperbolic GRU NQS in the form of the complex NQS wavefunction Eq.(45) described in 4.2. The architecture of these neural networks consist of a Euclidean/hyperbolic RNN unit and two Dense layers with 2 units each, one with the Softmax activation and one with the Softsign activation. Furthermore, making use of the observation from [9] that reported an improvement in the final results when the Marshall sign [47] is used, the wavefunction can be written as
$\Psi(\vec{\sigma})=(-1)^{M_{A}(\vec{\sigma})}\tilde{\Psi}(\vec{\sigma})$ (54)
where $M_{A}(\vec{\sigma})=\sum_{i\in A}{\sigma}_{i}$, and $\tilde{\Psi}({\sigma})$ is the positive amplitude of the wavefunction, with $A$ comprising the sites belonging to the sublattice of all even or all odd sites in the lattice. The factor $(-1)^{M_{A}(\vec{\sigma})}$ is the Marshall sign of the wavefunction, and Eq.(54) holds true when $J_{2}=0$. When $J_{2}\neq 0$, Eq.(54) is no longer exact but still holds approximately [7].
In all VMC experiments, the same Euclidean GRU variant, eGRU-75-s50, with the RNN hidden dimension $d_{h}=75$ and 17854 trainable parameters [^14] As a passing remark, we note that in the work [9], the authors used a much larger GRU-based model, comprising three layers of GRU, each with $d_{h}=100$ to approximate the ground state energy of the 1D $J_{1}J_{2}$ model with $N=100$ spins. Their goal was to establish the performance of GRU-based neural network as an effective NQS. Here we are interested in establishing the performance of the hyperbolic version of the GRU-based NQS and not its Euclidean counterpart. This, and the fact that our 1D $J_{1}J_{2}$ system has a smaller number of spins $N$ ($N=50$ instead of 100), motivates our choice of using a smaller Euclidean GRU-based NQS. is used to establish a benchmark performance against which we compare the performances of different variants of hyperbolic GRU. In particular, for each $J_{2}$ value, three variants of hyperbolic GRU with the RNN hidden dimension of $d_{h}=60,70,75$ (hGRU-60-s50, hGRU-70-s50, hGRU-75-s50) were considered and used to run the VMC experiments, and the best result among these variants [^15] Sometimes a variant with smaller $d_{h}$ hidden units outperforms one with larger $d_{h}$, as is the case for $J_{2}=0.0,0.5$ when hGRU-60-s50 and hGRU-70-s50 outperform hGRU-75-s50, which ran into the problem of overfitting during the training process. is selected to compare with the benchmark result provided by the Euclidean GRU. The list of all Euclidean/hyperbolic GRU variants is listed in Table 7.
@@
Aside from the fact that the wavefunction in this case is complex instead of real, the other training details (the type of optimizers used, the training method in which only the best model is saved, and the hardware platform) remain exactly the same as described in the 1D TFIM case. The number of training epochs are 450 for $J_{2}=0.0,0.2$ and 400 for $J_{2}=0.5,0.8$. As is the case in the 1D TFIM setting, hyperbolic GRU takes 5-7 times longer to complete the training loop than Euclidean GRU due to their more involved mathematical construction [^16] For $J_{2}=0$, Euclidean GRU takes around 2.7 hours to complete the training, while for $J_{2}\neq 0$, Euclidean GRU takes around 4 hours to complete the training on an average laptop with no GPU.
@@ 5.3.2 Results
In Table 8 we list the results recorded during the training process and in Table 9 we list the results obtained by performing the inference process in which the trained models are used to generate new samples to calculate new NQS wavefunction. Fig.8 graphically represent the data in both Tables 8, 9 and show the comparison of the performances between Euclidean and hyperbolic GRU for all four values of $J_{2}$. The convergence curve plots for the mean energy and energy variance recorded during the training process are shown in Fig.15 and Fig.16, respectively, in Appendix A.3. Unlike the smooth convergence curves of the 1D TFIM case, the training in this case is less stable with multiple kinks in the energy convergence curves (see Fig.15) as well as multiple spikes in the variance convergence curves (Fig.16).
@@
Table 8: VMC training results for the 1D $J_{1}J_{2}$ system with the Hamiltonian Eq.(53) with $N=50$, $J_{1}=1$ and $J_{2}=0.0,0.2,0.5,0.8$. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). The best result at each $J_{2}$ value is noted in bold.
Ansatz $J_{2}=0.0$ $J_{2}=0.2$ $J_{2}=0.5$ $J_{2}=0.8$
Euclidean GRU -21.2575 -17.7340 -17.3381 -19.3249
(0.59) (2.05) (0.63) (2.02)
Hyperbolic GRU -20.3342 -19.9232 -18.4635 -20.1790
(0.75) (0.58) (0.58) (1.05)
Exact (DMRG) -21.9721 -20.3150 -18.7500 -20.9842
@@
Table 9: VMC inference results for the 1D $J_{1}J_{2}$ system with the Hamiltonian Eq.(53) with $N=50$, $J_{1}=1$ and $J_{2}=0.0,0.2,0.5,0.8$. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). The best result at each $J_{2}$ value is noted in bold.
@@
Figure 8: Comparisons of the performances of Euclidean and hyperbolic GRU NQS ansatzes listed in Table 7 from VMC runs of 1D $J_{1}J_{2}$ model with different $J_{2}$ values (clockwise from top: $J_{2}$ =0.0, 0.2, 0.8, 0.5) for $N=50$ spins. In each subfigure, the mean energy and standard error from training and inference for each NQS ansatz is shown as a dot with error bar. The specific variants of Euclidean/hyperbolic GRU used in each of the four cases are listed in Table 7. In this plot, we do not specify the exact variants, but label them simply as Euclidean GRU or hyperbolic GRU to emphasize the geometry of the ansatzes.
@@
From Table 8, Table 9 and Fig.8, we note that in both training and inference, hyperbolic GRU outperform Euclidean GRU on 3 out of 4 cases. In particular, these 3 cases all correspond to nonzero values of $J_{2}$, i.e. when $J_{2}=0.2,0.5,0.8$, in which the next nearest neighbor interactions are taken into account. When $J_{2}=0.0$ (only nearest interactions - no next nearest neighbor interations), Euclidean GRU outperform hyperbolic GRU (in both training and inference). This result is again very interesting in the sense that it might be again indicative of the connection between the geometry of the NQS ansatz used and the structure of the Hamiltonian system under study, similar to the 2D TFIM setting above in which we observe a clear trend of outperformance by 1D hyperbolic GRU compared to 1D Euclidean GRU when there is a hierarchy in the Hamiltonian interaction structure, comprising the first and $N^{th}$ nearest neightbor interactions, introduced by the artificial rearrangement of the 1D spin chain to mimic the 2D spin lattice. In this case, when second nearest neighbor interactions are switched on ($J_{2}>0$), there is also a hierarchy going from the first nearest to the second nearest interactions, and it is our hypothesis that it is the decisive factor that leads to the outperformance of the hyperbolic GRU-based NQS over its Euclidean counterpart. It is unclear to us at the moment, although it is within our expectation, whether this would still be the case in two dimenions, i.e. whether hyperbolic GRU NQS carries on outperforming Euclidean GRU NQS in 2D $J_{1}J_{2}$ models with different 2D lattice geometries (such as square, triangle, kagome, etc.). In order to perform the VMC experiments for 2D $J_{1}J_{2}$ systems, one first needs to generalize the 1D hyperbolic GRU part of the NQS to two dimensions, similar to what was done in [9] in generalizing the 1D Euclidean RNN cell to its 2D version in Eq.(52).
@@ 5.4.1 Experiment details
  • In the first set of experiments, all four ($J_{2},J_{3}$) cases listed above are considered, withnthe Euclidean/hyperbolic GRU NQS ansatzes used having the same RNN hidden dimension of $d_{h}=50$ (see Table 10 for the variant names and their numbers of parameters).
@@
  • In the second set of experiments, we only considered $(J_{2},J_{3})$ = (0.2, 0.5), (0.5, 0.2), with the Euclidean GRU ansatz having the RNN hidden dimension $d_{h}=60$ and the hyperbolic GRU ansatz having $d_{h}=55$ (see Table 11).
@@
    Table 11: The variants of Euclidean and hyperbolic GRU used as NQS ansatzes to run the second set of VMC experiments involving the 1D $J_{1}J_{2}J_{3}$ Hamiltonian Eq.(55) with $N=30$, $J_{1}=1$ and $(J_{2},J_{3})=(0.2,0.5),(0.5,0.2)$. Only one variant of Euclidean GRU, eGRU-60-s50 with $d_{h}=60$ was used. Three variants of hyperbolic GRU with $d_{h}=53,55,57$ were used in the experiment, but we only list the best performing variant in this Table. The number of parameters of each NQS is listed in brackets under their names.
In the first set of VMC experiments, we only use one variant each for either Euclidean (eGRU-50-s50 with $d_{h}=50$) or hyperbolic GRU (hGRU-50-s50 with $d_{h}=50$) and fix the number of training epochs to 280 in all cases. In the second set of VMC experiments, we explore the performances of hyperbolic GRU variants with $d_{h}=53,55,57$ to choose the best one among them to compare with the Euclidean GRU benchmark, eGRU-60-s50 with $d_{h}=60$, while extending the number of training epochs to 350 for $(J_{2},J_{3})=(0.2,0.5)$ and 500 for $(J_{2},J_{3})=(0.5,0.2)$. Similar to previous cases, in each experiment, hyperbolic GRU can take up to around 5 times longer to complete the training (on the same hardware) than Euclidean GRU with the same $d_{h}$171717while hyperbolic GRU with smaller $d_{h}$ also takes longer to train, up to around 2-3 times, than Euclidean GRU with larger $d_{h}$. All other training details (optimizer settings, saving-of-best-model-only method) remain the same as in the 1D $J_{1}J_{2}$ case and will not be repeated here.
@@ 5.4.2 Results
  • For the first set of VMC experiments involving $(J_{2},J_{3})$ = (0.0, 0.5), (0.2, 0.2), (0.2, 0.5), (0.5, 0.2) with the NQS ansatzes listed in Table 10, the training and inference results are listed in Table 12 and Table 13, respectively. Fig.9 graphically shows the data in both these Tables and offers a comparison between the performances of the 1D Euclidean and 1D hyperbolic GRU. The convergence curve plots for the mean energy and energy variance recorded during the training process are shown in Fig.17 and Fig.18, respectively, in Appendix A.4.
@@
    Hyperbolic GRU -15.1557 -12.2377 -12.9018 -10.954
    (1.690) (0.659) (0.4532) (0.3374)
    Exact (DMRG) -15.89028 -12.9430 -14.64083 -11.52874
    Table 12: VMC training results (from the first set of experiments with the NQS ansatzes listed in Table 10) for the 1D $J_{1}J_{2}J_{3}$ system with the Hamiltonian Eq.(55) with $N=30$, $J_{1}=1$, and $(J_{2},J_{3})=(0.0,0.2),(0.2,0.2),(0.2,0.5),(0.5,0.2)$. These results are recorded directly from the training and correspond to the parameters of the best saved model. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). The best result at each $(J_{2},J_{3})$ case is noted in bold.
    Euclidean GRU -11.3272 -10.7929 -9.9025 -9.74123
    (0.002) (0.908) (0.001) (1.547)
    Hyperbolic GRU -15.2398 -12.2060 -12.$6$10.7659
@@
    Table 13: VMC inference results (from the first set of experiments with the NQS ansatzes listed in Table 10) for the 1D $J_{1}J_{2}J_{3}$ system with the Hamiltonian Eq.(55) with $N=30$, $J_{1}=1$, and $(J_{2},J_{3})=(0.0,0.2),(0.2,0.2),(0.2,0.5),(0.5,0.2)$. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). The best result at each $(J_{2},J_{3})$ case is noted in bold.
    Figure 9: Comparisons of the performances of Euclidean and hyperbolic GRU NQS ansatzes listed in Table 10 from VMC runs of 1D $J_{1}J_{2}J_{3}$ model with different $(J_{2},J_{3})$ values - clockwise from top: (0.0, 0.5), (0.2, 0.2), (0.5, 0.2), (0.2, 0.5) - for $N=30$ spins. In each subfigure, the mean energy and standard error from training and inference for each NQS ansatz is shown as a dot with an error bar. As is the $J_{1}J_{2}$ case, we do not specify the exact variants (listed in Table 10) in this plot, but choosing instead to label them as ‘Euclidean GRU’ and ‘Hyperbolic GRU’ to emphasize the geometry of the NQS ansatz.
    From Table 12, Table 13 and Fig.9, we note that in both training and inference, hyperbolic GRU definitively outperformed Euclidean GRU (by a large margin) in all 4 out of 4 cases. The energy convergence curves (see Fig.17) corresponding to the hyperbolic GRU ansatz reach convergence at values well below that reached by the curves corresponding to the Euclidean GRU ansatz, while the variance convergence curves of hyperbolic GRU (see Fig.18) show smaller ranges of fluctuations than Euclidean GRU’s.
  • For the second set of VMC experiments involving only ($J_{2},J_{3})=(0.2,0.5),(0.5,0.2)$ with the NQS ansatzes listed in Table 11, the training and inference results are included in Table 14 and Table 15. The comparison of the performances of hyperbolic and Euclidean GRU ansatzes is shown in Fig.10. The convergence curve plots of the second set of VMC experiments for the mean energy and energy variance recorded during the training process are shown in Fig.19 and Fig.20, respectively, in Appendix A.4.
@@
    Table 14: VMC training results (from the second set of experiments with the NQS ansatzes listed in Table 11) for the 1D $J_{1}J_{2}J_{3}$ system with the Hamiltonian Eq.(55) with $N=30$, $J_{1}=1$, and $(J_{2},J_{3})=(0.2,0.5),(0.5,0.2)$. These results are recorded directly from the training and correspond to the parameters of the best saved model. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). The best result at each $(J_{2},J_{3})$ case is noted in bold.
    Ansatz $(J_{2},J_{3})=(0.2,0.5)$ $(J_{2},J_{3})=(0.5,0.2)$
    Euclidean GRU -14.6828 -11.2061
    (1.000) (0.403)
    Hyperbolic GRU -14.4243 -11.3070
    (0.640) (0.065)
    Exact (DMRG) -14.6408 -11.5287
@@
    Table 15: VMC inference results (from the second set of experiments with the NQS ansatzes listed in Table 11) for the 1D $J_{1}J_{2}J_{3}$ system with the Hamiltonian Eq.(55) with $N=30$, $J_{1}=1$, and $(J_{2},J_{3})=(0.2,0.5),(0.5,0.2)$. For each ansatz, the mean $E$ is listed first, together with the variance (in brackets). The best result at each $(J_{2},J_{3})$ case is noted in bold.
@@
    Figure 10: Comparisons of the performances of Euclidean and hyperbolic GRU NQS ansatzes listed in Table 11 from VMC runs of 1D $J_{1}J_{2}J_{3}$ model with different $(J_{2},J_{3})$ values - from left to right: (0.2, 0.5), (0.5, 0.2) - for $N=30$ spins. In each subfigure, the mean energy and standard error from training and inference for each NQS ansatz is shown as a dot with an error bar. As before, we do not specify the exact variants (listed in Table 11) in this plot, but choosing instead to label them as ‘Euclidean GRU’ and ‘Hyperbolic GRU’ to emphasize the geometry of the NQS ansatz.
@@
    From Table 14, Table 15 and Fig.10, for $(J_{2},J_{3})$ = (0.2, 0.5), the hyperbolic GRU ansatz hGRU-55-s50 outperforms the Euclidean GRU ansatz eGRU-60-s50 in the training process, as is evidence from the closer mean energy value obtained during training, as well as from the energy convergence curve shown in the first subfigure of Fig.19, which shows the curve corresponding to hGRU-55-s50 reaching convergence faster with a narrower fluctuation range than the curve corresponding to eGRU-60-s50. In this case, it must be noted that the Euclidean GRU ansatz used, eGRU-60-s50, has around 20 percent more parameters (11584 versus 9794) than the hyperbolic GRU ansatz, hGRU-55-s50. For $(J_{2},J_{3})$ = (0.5, 0.2), the hyperbolic GRU ansatz hGRU-57-s50 outperforms the Euclidean GRU ansatz eGRU-60-s50 in both the training and inference processes. In this case, hyperbolic GRU has 10 percent fewer parameters than Euclidean GRU (10492 versus 11584).
@@
From the two sets of VMC experiments for the 1D $J_{1}J_{2}J_{3}$ model, we observe that the obtained results showing that hyperbolic GRU almost always outperform Euclidean GRU are in agreement with the ones in the 1D $J_{1}J_{2}$ case in which hyperbolic GRU outperfomed its Euclidean version in all cases when $J_{2}\neq 0$, i.e. when the second neighbor interaction is nonvanishing and the Hamiltonian exhibits a hierarchy in the interaction structure comprised of the first and second neighbor interactions. In all four $J_{1}J_{2}J_{3}$ settings where there exist an even more pronounced hierarchy in the interaction structure of the Hamiltonian that comprises the first, second and third nearest neighbor interactions, the clear outperformance of the hyperbolic GRU ansatz reinforces the hypothesis that the hyperbolic geometry of this neural network plays a determining role in handling the type of Hamiltonians with hierarchical structures such as the 1D $J_{1}J_{2}$, $J_{1}J_{2}J_{3}$ systems. As already noted in the 1D $J_{1}J_{2}$ case, an interesting task would be to verify whether hyperbolic GRU carries on outperforming Euclidean GRU for the 2D $J_{1}J_{2}J_{3}$ models on different lattice geometries, but this will require the generalization of the 1D hyperbolic GRU network to a 2D version, similar to the generalization of the 1D Euclidean RNN cell to the 2D version in Eq.52 as done in [9].

### v3

@@ 5.3.1 Experiment details
For the 1D $J_{1}J_{2}$ system with $J_{1}=1.0$ and $J_{2}=0.0,0.2,0.5,0.8$, we carried out several VMC experiments using Euclidean and hyperbolic GRU NQS in the form of the complex NQS wavefunction Eq.(45) described in 4.2. The architecture of these neural networks consist of a Euclidean/hyperbolic RNN unit and two Dense layers with 2 units each, one with the Softmax activation and one with the Softsign activation.
In all VMC experiments, the same Euclidean GRU variant, eGRU-75-s50, with the RNN hidden dimension $d_{h}=75$ and 17854 trainable parameters [^17] As a passing remark, we note that in the work [9], the authors used a much larger GRU-based model, comprising three layers of GRU, each with $d_{h}=100$ to approximate the ground state energy of the 1D $J_{1}J_{2}$ model with $N=100$ spins. Their goal was to establish the performance of GRU-based neural network as an effective NQS. Here we are interested in establishing the performance of the hyperbolic version of the GRU-based NQS and not its Euclidean counterpart. This, and the fact that our 1D $J_{1}J_{2}$ system has a smaller number of spins NN ($N=50$ instead of 100), motivates our choice of using a smaller Euclidean GRU-based NQS. is used to establish a benchmark performance against which we compare the performances of different variants of hyperbolic GRU. In particular, for each $J_{2}$ value, three variants of hyperbolic GRU with the RNN hidden dimension of $d_{h}=60,70,75$ (hGRU-60-s50, hGRU-70-s50, hGRU-75-s50) were considered and used to run the VMC experiments, and the best result among these variants [^18] Sometimes a variant with smaller $d_{h}$ hidden units outperforms one with larger $d_{h}$, as is the case for $J_{2}=0.0,0.5$ when hGRU-60-s50 and hGRU-70-s50 outperform hGRU-75-s50, which ran into the problem of overfitting during the training process. is selected to compare with the benchmark result provided by the Euclidean GRU. The list of all Euclidean/hyperbolic GRU variants is listed in Table 5.
@@
Aside from the fact that the wavefunction in this case is complex instead of real, the other training details remain exactly the same as described previously. The number of training epochs are 450 for all experiments. Similar to previous cases, hyperbolic GRU takes 5-7 times longer to complete the training loop than Euclidean GRU due to their more involved mathematical construction [^19] For $J_{2}=0$, Euclidean GRU takes around 2.7 hours to complete the training, while for $J_{2}\neq 0$, Euclidean GRU takes around 4 hours to complete the training on an average laptop with no GPU.
@@ 5.3.2 Results
In Table 6 we list the results obtained by performing the inference process in which the trained models are used to generate $10^{4}$ new samples to calculate new NQS wavefunction. Fig.6 graphically represents the data in Table 6 and shows the comparison of the performances between Euclidean and hyperbolic GRU NQS for all four values of $J_{2}$. In this case, unlike the TFIM cases, we note that the use of gradient clipping (both by value and by global norm [^20] We performed the experiments using gradient clipping by both value and norm, and chose the best results among these.) was necessary [^21] The use of gradient clipping was necessary in almost all cases, with the exception of hGRU-70-s50 in the case of $J_{2}=0.5$. No gradient clipping was used during the training of hyperbolic GRU in that case (when $J_{2}=0.5$), as the energy curve for hGRU-70-s50 showed no signs of any kinks or jumps. to achieve stability (to avoid large jumps or kinks in the energy convergence curve) during the training process. The convergence curve plots for the mean energy recorded during the training process with gradient clipping are shown in Fig.15 in Appendix B.3.
For completeness, we also included, in Appendix B.3 the results obtained by training the NQS without using gradient clipping (see Table 11 and Fig.13 for the full results, see also Fig.14 for the energy convergence graphs). Furthermore, Fig.16 and Fig.17 show the energy convergence curves obtained during training with and without gradient clipping for Euclidean GRU and hyperbolic GRU, respectively.
@@
Table 6: VMC inference results (using 10000 samples) for the 1D $J_{1}J_{2}$ system with the Hamiltonian Eq.(53) with $N=50$, $J_{1}=1$ and $J_{2}=0.0,0.2,0.5,0.8$. For each ansatz, the mean $E$ is listed first, together with the corresponding standard error (in brackets). The best result at each $J_{2}$ value is noted in bold.
@@
Figure 6: Comparisons of the performances of Euclidean and hyperbolic GRU NQS ansatzes listed in Table 5 from VMC runs of 1D $J_{1}J_{2}$ model with different $J_{2}$ values (clockwise from top: $J_{2}$ =0.0, 0.2, 0.8, 0.5) for $N=50$ spins. In each subfigure, the mean energy for each NQS ansatz is shown as a dot with an error bar (representing the standard error). The specific variants of Euclidean/hyperbolic GRU used in each of the four cases are listed in Table 5. In this plot, we do not specify the exact variants, but label them simply as Euclidean GRU or hyperbolic GRU to emphasize the geometry of the ansatzes.
From Table 6 and Fig.6, we note that hyperbolic GRU outperformed Euclidean GRU in three out of four cases. In particular, these three cases all correspond to nonzero values of $J_{2}$, i.e. when $J_{2}=0.2,0.5,0.8$, in which the next nearest neighbor interactions $J_{2}$ are taken into account. When $J_{2}=0.2,0.8$, hyperbolic GRU definitively outperformed Euclidean GRU, while for $J_{2}=0.5$, the two architectures showed almost comparable performances, with the hyperbolic GRU emerging as the marginally better performer. On the other hand, when $J_{2}=0.0$ (only nearest interactions $J_{1}$ - no next nearest neighbor interations $J_{2}$), Euclidean GRU outperformed hyperbolic GRU. It is noteworthy to point out that without gradient clipping, for $J_{2}\neq 0$, Euclidean GRU performed much worse than hyperbolic GRU (as is evident from Table 11 and Fig.13). The use of gradient clipping eliminated the kinks/jumps in the energy convergence curves and improved the performances of the Euclidean GRU NQS significantly in the cases of $J_{2}=0.2$ and $J_{2}=0.5$ (see Fig.16 for a comparison of the energy convergence curves with and without gradient clipping). On the other hand, the use of gradient clipping only improved the performances of hyperbolic GRU NQS marginally, with the most notable improvements seen in the cases of $J_{2}=0.0$ and $J_{2}=0.8$ (see Fig.17). While gradient clipping improved the overall performances of both Euclidean and hyperbolic GRU (and much more so in the case of Euclidean GRU than hyperbolic GRU), the latter still ended up outperforming the former.
@@
This result is again very interesting in the sense that it might be again indicative of the connection between the geometry of the NQS ansatz used and the structure of the Hamiltonian system under study, similar to the 2D TFIM setting above in which we observe a clear trend of outperformance by 1D hyperbolic GRU compared to 1D Euclidean GRU when there is a hierarchy in the Hamiltonian interaction structure, comprising the first and $N^{th}$ nearest neighbor interactions, introduced by the artificial rearrangement of the 1D spin chain to mimic the 2D spin lattice. In this case, when second nearest neighbor interactions are switched on ($J_{2}>0$), there is also a hierarchy going from the first nearest to the second nearest interactions, and it is our hypothesis that it is the decisive factor that leads to the outperformance of the hyperbolic GRU-based NQS over its Euclidean counterpart. It is unclear to us at the moment, although it is within our expectation, whether this would still be the case in two dimenions, i.e. whether hyperbolic GRU NQS carries on outperforming Euclidean GRU NQS in 2D $J_{1}J_{2}$ models with different 2D lattice geometries (such as square, triangle, kagome, etc.). In order to perform the VMC experiments for 2D $J_{1}J_{2}$ systems, one first needs to generalize the 1D hyperbolic GRU part of the NQS to two dimensions, similar to what was done in [9] in generalizing the 1D Euclidean RNN cell to its 2D version in Eq.(55).
@@ 5.4.1 Experiment details
  • In the first set of experiments, all four ($J_{2},J_{3}$) cases listed above are considered, with the Euclidean/hyperbolic GRU NQS ansatzes used having the same RNN hidden dimension of $d_{h}=50$ (see Table 7 for the variant names and their numbers of parameters).
@@
  • In the second set of experiments, we only considered $(J_{2},J_{3})$ = (0.2, 0.5), (0.5, 0.2), with the Euclidean GRU ansatz having the RNN hidden dimension $d_{h}=60$. We experimented with various hyperbolic GRU ansatzes, each one having a different $d_{h}$ where $53\leq d_{h}\leq 58$ in order to choose the best performing ansatz to compare the Euclidean NQS (see Table 8).
@@
    Table 8: The variants of Euclidean and hyperbolic GRU used as NQS ansatzes to run the second set of VMC experiments involving the 1D $J_{1}J_{2}J_{3}$ Hamiltonian Eq.(54) with $N=30$, $J_{1}=1$ and $(J_{2},J_{3})=(0.2,0.5),(0.5,0.2)$. Only one variant of Euclidean GRU, eGRU-60-s50 with $d_{h}=60$ was used. Different variants of hyperbolic GRU with $53\leq d_{h}\leq 58$ were used in the experiment, but we only list the best performing variant in this Table. The number of parameters of each NQS is listed in brackets under their names.
In the first set of VMC experiments, we only used one variant each for either Euclidean (eGRU-50-s50 with $d_{h}=50$) or hyperbolic GRU (hGRU-50-s50 with $d_{h}=50$) and fixed the number of training epochs to 280 in all cases. In the second set of VMC experiments, we explored the performances of hyperbolic GRU variants with $53\leq d_{h}\leq 58$ to choose the best one among them to compare with the Euclidean GRU benchmark, eGRU-60-s50 with $d_{h}=60$, while extending the number of training epochs to 450 for $(J_{2},J_{3})=(0.2,0.5)$ and 500 for $(J_{2},J_{3})=(0.5,0.2)$. Similar to previous cases, in each experiment, hyperbolic GRU can take up to around 5 times longer to complete the training (on the same hardware) than Euclidean GRU with the same $d_{h}$222222while hyperbolic GRU with smaller $d_{h}$ also takes longer to train, up to around 2-3 times, than Euclidean GRU with larger $d_{h}$. All other training details (optimizer settings, saving-of-best-model-only method) remain the same as in the 1D $J_{1}J_{2}$ case and will not be repeated here.
@@ 5.4.2 Results
  • For the first set of VMC experiments involving $(J_{2},J_{3})$ = (0.0, 0.5), (0.2, 0.2), (0.2, 0.5), (0.5, 0.2) with the NQS ansatzes listed in Table 7, the inference results using $10^{4}$ samples are listed in Table 9, respectively. Fig.7 graphically shows the data in both these Tables and offers a comparison between the performances of the 1D Euclidean and 1D hyperbolic GRU. The convergence curve plots for the mean energy recorded during the training process are shown in Fig.18 in Appendix B.4.
@@
    Hyperbolic GRU -15.0637 -11.9804 -12.7448 -10.7973
@@
    Table 9: VMC inference results (from the first set of experiments with the NQS ansatzes listed in Table 7) for the 1D $J_{1}J_{2}J_{3}$ system with the Hamiltonian Eq.(54) with $N=30$, $J_{1}=1$, and $(J_{2},J_{3})=(0.0,0.2),(0.2,0.2),(0.2,0.5),(0.5,0.2)$. For each ansatz, the mean $E$ is listed first, together with the standard error below (in brackets). The best result at each $(J_{2},J_{3})$ case is noted in bold.
    Figure 7: Comparisons of the performances of Euclidean and hyperbolic GRU NQS ansatzes listed in Table 7 from VMC runs of 1D $J_{1}J_{2}J_{3}$ model with different $(J_{2},J_{3})$ values - clockwise from top: (0.0, 0.5), (0.2, 0.2), (0.5, 0.2), (0.2, 0.5) - for $N=30$ spins. In each subfigure, the mean energy for each NQS ansatz is shown as a dot with an error bar (representing the standard error). As is the $J_{1}J_{2}$ case, we do not specify the exact variants (listed in Table 7) in this plot, but choosing instead to label them as ‘Euclidean GRU’ and ‘Hyperbolic GRU’ to emphasize the geometry of the NQS ansatz.
    From Table 9 and Fig.7, we note that hyperbolic GRU definitively outperformed Euclidean GRU (by a large margin) in all 4 out of 4 cases. The energy convergence curves (see Fig.18) corresponding to the hyperbolic GRU ansatz reach convergence at values well below that reached by the curves corresponding to the Euclidean GRU ansatz.
  • For the second set of VMC experiments involving only ($J_{2},J_{3})=(0.2,0.5),(0.5,0.2)$ with the NQS ansatzes listed in Table 8, the inference results are included in Table 10. The comparison of the performances of hyperbolic and Euclidean GRU ansatzes is shown in Fig.8. The convergence curve plots of the second set of VMC experiments for the mean energy recorded during the training process are shown in Fig.19 in Appendix B.4.
@@
    Table 10: VMC inference results (from the second set of experiments with the NQS ansatzes listed in Table 8) for the 1D $J_{1}J_{2}J_{3}$ system with the Hamiltonian Eq.(54) with $N=30$, $J_{1}=1$, and $(J_{2},J_{3})=(0.2,0.5),(0.5,0.2)$. For each ansatz, the mean $E$ is listed first, together with the corresponding standard error directly below (in brackets). The best result at each $(J_{2},J_{3})$ case is noted in bold.
@@
    Figure 8: Comparisons of the performances of Euclidean and hyperbolic GRU NQS ansatzes listed in Table 8 from VMC runs of 1D $J_{1}J_{2}J_{3}$ model with different $(J_{2},J_{3})$ values - from left to right: (0.2, 0.5), (0.5, 0.2) - for $N=30$ spins. In each subfigure, the mean energy for each NQS ansatz is shown as a dot with an error bar (representing the standard error). As before, we do not specify the exact variants (listed in Table 8) in this plot, but choosing instead to label them as ‘Euclidean GRU’ and ‘Hyperbolic GRU’ to emphasize the geometry of the NQS ansatz.
    1. 1.
@@
      For $(J_{2},J_{3})$ = (0.2, 0.5), the hyperbolic GRU ansatz hGRU-55-s50, with 9794 parameters, shows an overall comparable performance to the Euclidean GRU ansatz eGRU-60-s50 with 11584 parameters. Although the larger Euclidean GRU ansatz emerges as the better performer by a slight margin, it also has 20 percent more parameters than the hyperbolic GRU network. Furthermore, it is noteworthy to point out that in the training process, hGRU-55-s50 performed better than eGRU-60-s50, as is evident from the energy convergence curve shown in the first subfigure of Fig.19, which shows the curve corresponding to hGRU-55-s50 reaching convergence faster with a narrower fluctuation range than the curve corresponding to eGRU-60-s50.
    2. 2.
      For $(J_{2},J_{3})$ = (0.5, 0.2), the hyperbolic GRU ansatz hGRU-57-s50 outperforms the Euclidean GRU ansatz eGRU-60-s50. In this case, hyperbolic GRU has 10 percent fewer parameters than Euclidean GRU (10492 versus 11584).
@@
From the two sets of VMC experiments for the 1D $J_{1}J_{2}J_{3}$ model, we observe that the obtained results showing that hyperbolic GRU almost always outperformed Euclidean GRU are in agreement with the ones in the 1D $J_{1}J_{2}$ case in which hyperbolic GRU outperfomed its Euclidean version in all cases when $J_{2}\neq 0$, i.e. when the second neighbor interaction is nonvanishing and the Hamiltonian exhibits a hierarchy in the interaction structure comprised of the first and second neighbor interactions. In all four $J_{1}J_{2}J_{3}$ settings where there exist an even more pronounced hierarchy in the interaction structure of the Hamiltonian that comprises the first, second and third nearest neighbor interactions, the clear outperformance of the hyperbolic GRU ansatz reinforces the hypothesis that the hyperbolic geometry of this neural network plays a determining role in handling the type of Hamiltonians with hierarchical structures such as the 1D $J_{1}J_{2}$, $J_{1}J_{2}J_{3}$ systems. As already noted in the 1D $J_{1}J_{2}$ case, an interesting task would be to verify whether hyperbolic GRU carries on outperforming Euclidean GRU for the 2D $J_{1}J_{2}J_{3}$ models on different lattice geometries, but this will require the generalization of the 1D hyperbolic GRU network to a 2D version, similar to the generalization of the 1D Euclidean RNN cell to the 2D version in Eq.55 as done in [9].